# Can Zero-Shot Commercial API’s Deliver Regulatory-Grade Clinical Text De-Identification?

### v1

@@ Abstract
We systematically assess the performance of three leading API-based de-identification systems—Azure Health Data Services, AWS Comprehend Medical, and OpenAI GPT-4o—against our de-identification systems on a ground truth dataset of 48 clinical documents annotated by medical experts. Our analysis, conducted at both entity-level and token-level, demonstrates that our solution, Healthcare NLP, achieves the highest accuracy, with a 96% F1-score in protected health information (PHI) detection, significantly outperforming Azure (91%), AWS (83%), and GPT-4o (79%). Beyond accuracy, Healthcare NLP is also the most cost-effective solution, reducing processing costs by over 80% compared to Azure and GPT-4o. Its fixed-cost local deployment model avoids the escalating per-request fees of cloud-based services, making it a scalable and economical choice. Our results underscore a critical limitation: zero-shot commercial APIs fail to meet the accuracy, adaptability, and cost-efficiency required for regulatory-grade clinical de-identification. Healthcare NLP’s superior performance, customization capabilities, and economic advantages position it as the more viable solution for healthcare organizations seeking compliance and scalability in clinical NLP workflows.
@@ 3.1 The Deidentification Solutions
3.1.1 Healthcare NLP & LLM Library
The Healthcare NLP library is a powerful component of Spark NLP platform [15], specifically designed to facilitate NLP tasks within the healthcare domain [16]. This library offers over 2,500 pre-trained models and pipelines tailored for medical data, enabling accurate information extraction, named entity recognition (NER) for clinical and medical concepts, and robust text analysis capabilities. Regularly updated with advanced algorithms, it helps healthcare professionals derive meaningful insights from unstructured medical data sources such as electronic health records, clinical notes, and biomedical literature.

### v2

@@ Abstract
We evaluate the performance of four leading solutions for de-identification of unstructured medical text - Azure Health Data Services, AWS Comprehend Medical, OpenAI GPT-4o, and John Snow Labs - on a ground truth dataset of 48 clinical documents annotated by medical experts. The analysis, conducted at both entity-level and token-level, suggests that John Snow Labs’ Medical Language Models solution achieves the highest accuracy, with a 96% F1-score in protected health information (PHI) detection, outperforming Azure (91%), AWS (83%), and GPT-4o (79%). John Snow Labs is not only the only solution which achieves regulatory-grade accuracy (surpassing that of human experts) but is also the most cost-effective solution: It is over 80% cheaper compared to Azure and GPT-4o, and is the only solution not priced by token. Its fixed-cost local deployment model avoids the escalating per-request fees of cloud-based services, making it a scalable and economical choice.
@@ 3.1 The Deidentification Solutions
3.1.1 John Snow Labs Healthcare NLP Library & Medical Language Models
The Healthcare NLP library by John Snow Labs is a powerful component of Spark NLP platform [15], specifically designed to facilitate NLP tasks within the healthcare domain [16]. This library offers over 2,500 pre-trained models and pipelines tailored for medical data, enabling accurate information extraction, named entity recognition (NER) for clinical and medical concepts, and robust text analysis capabilities. Regularly updated with advanced algorithms, it helps healthcare professionals derive meaningful insights from unstructured medical data sources such as electronic health records, clinical notes, and biomedical literature.